A pair of eyeglasses is designed to allow a person with a far-point distance of 2.44 m to read a road sign at a distance of 24.4 m .

a) Find the focal length required of these glasses if they are to be worn 1.50 cm from the eyes.

b) Find the focal length required of these glasses if they are to be worn 0.750 cm from the eyes.
Answer #1

a) as the lens formula,

1/f = 1/v + 1/u

now putting the values as given in question and applying sign conventions,

1/f = -1/(24.4-0.015) + 1/(2.44 - 0.015)

1/f = 0.3716

f = 2.691 m

b)

as the lens formula,

1/f = 1/v + 1/u

now putting the values as given in question and applying sign conventions,

1/f = -1/(24.4-0.0075) + 1/(2.44 - 0.0075)

1/f = 0.3701

f = 2.709 m